OpenStudy (anonymous):

In triangle ABC, the measure of angle C is eight times the measure of angle A, and the measure of angle B is 10° more than the measure of angle C. Find the measure of each angle of the triangle

OpenStudy (anonymous):

A+B+C=180 C=8A B=C+10=8A+10 So: A+(8A+10)+8A=180 17A=170 So A=10 then C=80 B=90